Jennifer had $30 to spend on herself. She spent 1/5 of the money on a sandwich, 1/6 for a ticket to a museum, and 1/2 of it on a book. How much money does Jennifer have left over?

 



1
Expert's answer
2020-10-05T17:45:32-0400

Jennifer spent "\frac{1}{5}+\frac{1}{6}+\frac{1}{2}=\frac{6}{30}+\frac{5}{30}+\frac{15}{30}=\frac{6+5+15}{30}=\frac{26}{30}" of money on sandwich, ticket and book and "1-\frac{26}{30}=\frac{30-26}{30}=\frac{4}{30}" of money Jennifer has left over.

It is "\frac{4}{30}\times \$ 30 =\$4".


Answer: $4.